OpenStudy (anonymous):

Identify the horizontal asymptote of f(x) = quantity 7 x plus 1 over quantity 2 x minus 9

OpenStudy (anonymous):

@camerondoherty

OpenStudy (imstuck):

There are rules about horizontal asymptotes. Here is your function:\[\frac{ 7x+1 }{ 2x-9 }\]Is that it?

OpenStudy (anonymous):

yes thats how it looks

OpenStudy (imstuck):

ok the rules for the horizontal asymptotes are as follows:

OpenStudy (imstuck):

If the power of the numerator's x is less than the power of the denominator's x, then the x axis is the hor.asymptote (y = 0 line). If the power of the numerator's x is equal to the power of the denominator's x, as yours is, the hor asymptote is the line y = a/b. Now your a is your coefficient on the x in the numerator (7), and your b is the coefficient on the x in the denominator (2). So your line is y = 7/2. That is your case. If the power of the x in the numerator is greater than the power of the x in the denominator, there is no hor asymptote.

OpenStudy (anonymous):

o ok thanks for the help appreciate it

OpenStudy (imstuck):

YW

OpenStudy (anonymous):

hey can you help me with one more

OpenStudy (anonymous):

Identify the oblique asymptote of f(x) = quantity x squared minus 4 x plus 8 over quantity x plus 2